The number of men attending Western Tech is 200 less than twice the number of women. The number of students at the school is 280. How man men attend? How many woman attend?
I am not sure how to come up with the correct Linear equation for this.. I tried the following
x= the # of men
2X-200=280
The answer I get is 240 but that is not what they were looking for... Can you help??

Answer by nerdybill(6950)

The number of men attending Western Tech is 200 less than twice the number of women. The number of students at the school is 280. How man men attend? How many woman attend?
.
Let w = number of women
then
2w-200 = number of men
.
w + 2w-200 = 280
3w-200 = 280
3w = 480
w = 160 (women)
.
Men:
2w-200 = 2(160)-200 = 320-200 = 120 (men)